# Scalable In-Context Q-Learning

### v2

@@ 1 Introduction
A longstanding goal of reinforcement learning (RL) is to learn from diverse experiences and generalize beyond its training environments, efficiently adapting to unseen situations, dynamics, or objectives (Ackley & Littman, 1989; Kirk et al., 2023). A promising avenue is in-context learning (Brown et al., 2020) that trains large-scale transformer models on massive datasets to achieve remarkable generalization capabilities, adapting to new tasks via prompt conditioning without any model updates (Wang et al., 2023; Li et al., 2024a). Accordingly, in-context RL (ICRL) seeks to extend this promise to decision domains and has seen rapid progress in recent years (Nikulin et al., 2025). Existing studies contain two typical branches: algorithm distillation (AD) (Laskin et al., 2023) and decision-pretrained transformer (DPT) (Lee et al., 2023), due to their simplicity and generality. They commonly employ cross-episode transitions as few-shot prompts and train transformer-based policies under supervised pretraining (Lin et al., 2024), followed by various improvements from model-based planning (Son et al., 2025), hierarchical decomposition (Huang et al., 2024), importance weighting (Dong et al., 2024), etc (Sinii et al., 2024; Tarasov et al., 2025; Dai et al., 2024).
@@
Though, significant challenges may emerge when extending the promise of in-context learning from (self-) supervised learning to RL, since RL involves more complex dynamics and temporal correlations (Silver et al., 2021). First, previous studies usually adopt a supervised pretraining paradigm, failing to go beyond imitating collected data (Yamagata et al., 2023). AD can require long-horizon context and inherit suboptimal behaviors due to the gradual update rule (Son et al., 2025). DPT relies on an oracle for optimal action labeling that can often be infeasible in practice (Tarasov et al., 2025). These limitations may hinder the efficient learning of optimal policies, especially when only suboptimal datasets are available. Second, substantial scope remains to advance the design of efficient prompts that can precisely encode RL task information. In language communities, prompts are concise, precise text instructions rich in semantic information, as text naturally conveys high-level concepts like objects (nouns) and actions (verbs) (Kakogeorgiou et al., 2022). However, ICRL approaches generally take raw transitions as prompts that have many more tokens than a sentence and can be highly redundant. Within transitions from offline datasets, the task information can be entangled with behavior policies, thus producing biased task inference at test time (Yuan & Lu, 2022). Hence, this type of prompt may be insufficient to precisely capture relevant information about decision tasks. The aforementioned limitations raise a key question: Can we design a scalable ICRL framework using lightweight prompts that precisely capture task-relevant information, while unleashing the core potential of fundamental reward maximization to learn from suboptimal data?
@@ 2 Related Work
The concept of agents adapting their behaviors within the context without model updates builds on earlier work in meta-RL, notably RL2 (Duan et al., 2016) that includes past interactions as inputs to an RNN-based policy. Recently, there has been a shift towards employing transformers to implement ICRL (Nikulin et al., 2025), driven by their proven ability to capture long-term dependencies and exhibit emergent in-context learning behaviors Brown et al. (2020). In online RL settings, AMAGO (Grigsby et al., 2024a; b) trains long-sequence transformers over entire rollouts with actor-critic learning to tackle in-context goal-conditioned problems, and (Lu et al., 2023) leverages the S4 (structured state space sequence) model’s ability to handle long-range sequences for ICRL tasks. In offline settings, two classical branches are AD (Laskin et al., 2023) and DPT (Lee et al., 2023). AD trains a causal transformer to autoregressively predict actions using preceding learning histories as context, while DPT predicts the optimal action based on a query state and a prompt of interaction transitions. Follow-up studies enhance in-context learning from different perspectives (Son et al., 2025; Huang et al., 2024; Sinii et al., 2024; Dai et al., 2024; Zisman et al., 2024). For example, IDT (Huang et al., 2024) designs a hierarchical decision structure, and DICP (Son et al., 2025) incorporates model-based planning with a learned dynamics model. These approaches generally adopt a supervised paradigm with raw transitions as prompts, while we harness dynamic programming for reward maximization and leverage world modeling to construct more efficient prompts.
DIT (Dong et al., 2024) and IC-IQL (Tarasov et al., 2025) are the most relevant to our work. DIT also uses a weighted maximum likelihood estimation loss to train the transformer policy, where the weights are directly calculated from observed rewards in collected datasets. In contrast, we learn separate in-context value functions for computing advantage weights, enabling more stable and sample-efficient policy extraction. IC-IQL integrates a Q-learning objective into AD. The policy and value networks are updated by optimizing only their heads, without propagating gradients to the transformer backbone. In contrast, we train the full multi-head transformer policy end-to-end, releasing the transformer’s scalability with a simplified pipeline. We also design a precise, lightweight prompt structure to overcome the limitation of AD algorithms that require long training histories as context. Empirical results in Sec. 4 demonstrate our superiority to these two baselines.
@@ 3.2 Scalable Model Architecture
World Modeling.
As shown in Figure 1-(a), we design a lightweight prompt structure capable of encoding precise information about decision tasks. The world model, representing environment dynamics $p(s^{\prime},r|s,a)$ (Hafner et al., 2025), fully characterizes the underlying task and remains invariant to behavior policies or the datasets collected. Inspired by this fundamental property of RL, we pretrain a generalized world model to acquire task-relevant information from the multi-task offline dataset. The world model contains a context encoder $E_{\phi}$ that abstracts recent $k$-step experiences $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$ into a task representation as $z^{i}_{t}\!=\!E_{\phi}(\eta_{t}^{i})$, and a dynamics decoder $D_{\varphi}$ that predict the instant reward and next state conditioned on the task representation as $\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$. Then, for training the in-context policy, we sample a short $h$-step trajectory randomly from dataset $\mathcal{D}^{i}$, and use the pretrained $E_{\phi}$ to transform raw transitions in this trajectory into compact task representations as
@@
Dynamic Programming.
Classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates, showcasing high stitching capacity that allows for not only imitating collected data but also achieving substantial improvement beyond it (Sutton et al., 1998). Inspired by this, we harness the stitching property to offer a promising avenue for unleashing ICRL’s potential toward explicit reward maximization. Following common practice in advanced offline RL algorithms (Kostrikov et al., 2022; Snell et al., 2023), we learn both a state value function $V$ and an action value function $Q$. As shown in Figure 1-(b), we design a prompt-based multi-head transformer architecture to simultaneously predict the optimal and value functions. Let $\theta$ denote parameters of the integrated transformer model. The model outputs the policy $\pi_{\theta}(a|s;\beta)$, the state value $V_{\theta}(s;\beta)$, and the action value $Q_{\theta}(s,a;\beta)$ using three separate heads, given the task prompt $\beta$ and corresponding query inputs ($s$ or $s,a$). Sec. 3.4 presents the detailed learning process.
@@ 3.3 World Model Pretraining for Prompt Construction
$\mathcal{L}(\phi,\varphi)=\mathbb{E}_{\eta_{t}^{i}\sim M^{i}}\!\left[\mathbb{E}_{z^{i}_{t}\sim E_{\phi}(\eta_{t}^{i})}\!\left[\lVert[r_{t},s_{t+1}]-D_{\varphi}(s_{t},a_{t};z_{t}^{i})\rVert^{2}\right]\right],~~~\forall~\text{task}~i.$ (3)
@@ 3.4 In-Context Q-Learning
where $L_{2}^{\omega}(u)=|\omega-\mathbbm{1}(u<0)|\cdot u^{2}$ is an asymmetric loss function with a expectile parameter $\omega\in(0.5,1)$. This form of expectile regression reduces the influence of Q<VQ\!<\!V predictions by a factor of $1\!-\!\omega$ while assigning more importance to Q>VQ\!>\!V predictions by a factor of $\omega$. In this way, we predict an upper-expectile of the temporal-difference target that approximates the maximum of $r(s_{t}^{i},a_{t}^{i})+\gamma Q_{\hat{\theta}}(s_{t+1}^{i},a_{t+1}^{i};\beta^{i})$ over actions $a_{t+1}^{i}$ constrained to the dataset actions. More details can be found in Implicit Q-Learning (Kostrikov et al., 2022).
In-Context Policy Extraction.
The value function learning procedure allows for stitching suboptimal trajectories to discover globally optimal behaviors. Then, we distill the in-context value functions into policy extraction with advantage-weighted regression (Peng et al., 2019), a supervised learning style that uses a simple and convergent maximum likelihood loss function as
@@ 3.5 Overall Optimization
$\mathcal{L}(\theta)=\mathcal{L}_{\pi}(\theta)+\mathcal{L}_{Q}(\theta)+\mathcal{L}_{V}(\theta).$ (7)
This design ensures that S-ICQL integrates lightweight prompt construction with a unified optimization pipeline, achieving efficient reward maximization and robust generalization across tasks.
@@ 4 Experiments
  • Can S-ICQL achieve consistent performance gain on few-shot generalization ability to unseen tasks compared to other strong baselines? (Sec. 4.1)
@@
Baselines. We compare to six competitive ICRL approaches and one offline meta-RL method, including: 1) IC-IQL (Tarasov et al., 2025), 2) DIT (Dong et al., 2024), 3) DICP (Son et al., 2025), 4) IDT (Huang et al., 2024), 5) AD (Laskin et al., 2023), 6) DPT (Lee et al., 2023), and 7) UNICORN (Li et al., 2024b). Details about these baselines are provided in Appendix E.
@@ 4.1 Main Results
The baselines typically work under the few-shot setting, since they rely on task prompts or warm-start data to infer task representations. We evaluate S-ICQL against these baselines in an aligned few-shot setting, where all methods utilize the same number of interaction trajectories for task inference. Figure 2 illustrates the evaluation return curves across various environments using Mixed datasets, and Table 1 summarizes the numerical results of converged performance. Across environments with varying reward functions and transition dynamics, S-ICQL consistently demonstrates superior data efficiency and higher asymptotic performance. In more complex environments such as Ant-Dir and HalfCheetah-Vel, the advantage of S-ICQL is more pronounced, highlighting its promising in-context learning abilities on challenging tasks. Moreover, S-ICQL generally exhibits lower variance during the testing phase, indicating both improved data efficiency and better training stability.
Notably, DPT yields suboptimal performance across most environments, underscoring the critical role of incorporating fundamental reward maximization into our framework. The significant improvement over DIT and IC-IQL also verifies that our method can provide a more efficient way for policy extraction with advantage-weighted regression or incorporating Q-learning objectives.
@@ 4.2 Ablation Study
Table 2 present ablation results across representative environments (refer to Appendix H for graphical results). First, ablating the world model causes a notable performance drop, especially in complex tasks like Ant-Dir, emphasizing its importance for precise in-context inference. Second, removing Q-learning reduces performance, underscoring its role in refining policies using suboptimal data. Integrating Q-learning enables S-ICQL to improve policies from noisy or inferior trajectories, enhancing its stitching ability to maximize rewards. Finally, removing both components results in the greatest decline, reducing the model to simple behavioral cloning that fails to generalize in complex tasks. Overall, the ablation results validate S-ICQL’s effectiveness in harnessing both world modeling and Q-learning for efficient reward maximization and generalization across tasks.
@@ 4.4 Stitching Capability
Figure 4: Comparison of best dataset returns with DPT and S-ICQL performances on training tasks.
We conducted an in-depth analysis to examine whether S-ICQL truly exhibits stitching capability. Specifically, we pretrained the model on offline datasets containing only suboptimal trajectories and evaluated it on the corresponding training tasks. Figure 4 presents a comparison between the best returns in the offline datasets (denoted as OPT for optimal dataset return)and the performance achieved by DPT, w/o_q (S-ICQL without Q-learning), and S-ICQL, including both the evaluation return curves and the final return values. As w/o_q and DPT follows the supervised pretraining paradigm, its performance never exceeds the dataset best, underscoring the limits of pure supervised methods. In contrast, S-ICQL achieves performance exceeding the best returns in the dataset on training tasks, providing strong evidence of genuine stitching capability by composing suboptimal trajectory segments into globally optimal policies.

### v3

@@ 1 Introduction
RL is a pivotal mechanism for training autonomous agents to solve complex tasks in interactive environments (Mnih et al., 2015), with expanding applications in frontier challenges such as fine-tuning language models (Yan et al., 2025; Hu et al., 2026; Zhan et al., 2026) and diffusion models (Black et al., 2024; Liu et al., 2026). A longstanding goal of RL is to learn from diverse experiences and generalize beyond its training environments, efficiently adapting to unseen situations, dynamics, or objectives (Ackley and Littman, 1989; Kirk et al., 2023; Hu et al., 2025). A promising avenue is in-context learning (Brown et al., 2020) that trains large-scale transformer models on massive datasets to achieve remarkable generalization capabilities, adapting to new tasks via prompt conditioning without any model updates (Wang et al., 2023a; Li et al., 2024a). Accordingly, in-context RL (ICRL) seeks to extend this promise to decision domains and has seen rapid progress in recent years (Nikulin et al., 2025). Existing studies contain two typical branches: algorithm distillation (AD) (Laskin et al., 2023) and decision-pretrained transformer (DPT) (Lee et al., 2023), due to their simplicity and generality. They commonly employ cross-episode transitions as few-shot prompts and train transformer-based policies under supervised pretraining (Lin et al., 2024), followed by various improvements from model-based planning (Son et al., 2025), hierarchical decomposition (Huang et al., 2024), importance weighting (Dong et al., 2025), etc (Sinii et al., 2024; Tarasov et al., 2025; Dai et al., 2024).
@@
Though, significant challenges may emerge when extending the promise of in-context learning from (self-) supervised learning to RL, since RL involves more complex dynamics and temporal correlations (Silver et al., 2021). First, previous studies usually adopt a supervised pretraining paradigm, failing to go beyond imitating collected data (Yamagata et al., 2023). AD can require long-horizon context and inherit suboptimal behaviors due to the gradual update rule (Son et al., 2025). DPT relies on an oracle for optimal action labeling that can often be infeasible in practice (Tarasov et al., 2025). These limitations may hinder the efficient learning of optimal policies, especially when only suboptimal datasets are available. Second, substantial scope remains to advance the design of efficient prompts that can precisely encode RL task information. In language communities, prompts are concise, precise text instructions rich in semantic information, as text naturally conveys high-level concepts like objects (nouns) and actions (verbs) (Kakogeorgiou et al., 2022). However, ICRL approaches generally take raw transitions as prompts that have many more tokens than a sentence and can be highly redundant. Within transitions from offline datasets, the task information can be entangled with behavior policies, thus producing biased task inference at test time (Yuan and Lu, 2022). Hence, this type of prompt may be insufficient to precisely capture relevant information about decision tasks. The aforementioned limitations raise a key question: Can we design a scalable ICRL framework using lightweight prompts that precisely capture task-relevant information, while unleashing the core potential of fundamental reward maximization to learn from suboptimal data?
@@ 2 Related Work
The concept of agents adapting their behaviors within the context without model updates builds on earlier work in meta-RL (Beck et al., 2025), such as the memory-based RL2 (Duan et al., 2016) and LLIRL (Wang et al., 2022; 2023b), the optimization-based MAML (Finn et al., 2017) and MACAW (Mitchell et al., 2021), and the context-based VariBAD (Zintgraf et al., 2021), Meta-DT (Wang et al., 2024), etc. (Li et al., 2024b; Zhang et al., 2025). Recently, there has been a shift towards employing transformers to implement ICRL (Moeini et al., 2025; Nikulin et al., 2025), driven by their proven ability to capture long-term dependencies and exhibit emergent in-context learning behaviors Brown et al. (2020). AMAGO (Grigsby et al., 2024a; b) trains long-sequence transformers over entire rollouts with actor-critic learning to tackle in-context goal-conditioned problems, and (Lu et al., 2023) leverages the S4 model’s ability to handle long-range sequences for ICRL tasks. In offline settings, two classical branches are AD (Laskin et al., 2023) and DPT (Lee et al., 2023). AD trains a causal transformer to autoregressively predict actions using preceding learning histories as context, while DPT predicts the optimal action based on a query state and a prompt of interaction transitions. Follow-up studies enhance in-context learning from different perspectives (Son et al., 2025; Huang et al., 2024; Dai et al., 2024; Zisman et al., 2024; Wu et al., 2025). For example, IDT (Huang et al., 2024) designs a hierarchical decision structure, and DICP (Son et al., 2025) incorporates model-based planning with a learned dynamics model. These approaches generally adopt a supervised paradigm with raw transitions as prompts, while we harness dynamic programming for reward maximization and leverage world modeling to construct more efficient prompts.
DIT (Dong et al., 2025) and IC-IQL (Tarasov et al., 2025) are the most relevant to our work. DIT also uses a weighted maximum likelihood estimation loss to train the transformer policy, where the weights are directly calculated from observed rewards in collected datasets. In contrast, we learn separate in-context value functions for computing advantage weights, enabling more stable and sample-efficient policy extraction. IC-IQL integrates a Q-learning objective into AD. The policy and value networks are updated by optimizing only their heads, without propagating gradients to the transformer backbone. In contrast, we train the full multi-head transformer policy end-to-end, releasing the transformer’s scalability with a simplified pipeline. We also design a precise, lightweight prompt structure to overcome the limitation of AD algorithms that require long training histories as context. Empirical results in Sec. 4 demonstrate our superiority to these two baselines.
@@ 3.2 Scalable Model Architecture
World Modeling. As shown in Figure 1-(a), we design a lightweight prompt structure capable of encoding precise information about decision tasks. The world model, representing environment dynamics $p(s^{\prime},r|s,a)$ (Hafner et al., 2025), fully characterizes the underlying task and remains invariant to behavior policies or the datasets collected. Inspired by this fundamental property of RL, we pretrain a generalized world model to acquire task-relevant information from the multi-task offline dataset. The world model contains a context encoder $E_{\phi}$ that abstracts recent $k$-step experiences $\eta_{t}^{i}\!=\!(s_{t-k},a_{t-k},r_{t-k},...,s_{t-1},a_{t-1},r_{t-1},s_{t},a_{t})^{i}$ into a task representation as $z^{i}_{t}\!=\!E_{\phi}(\eta_{t}^{i})$, and a dynamics decoder $D_{\varphi}$ that predict the instant reward and next state conditioned on the task representation as $\hat{r}_{t},\hat{s}_{t+1}]\!=\!D_{\varphi}(s_{t},a_{t};z_{t}^{i})$. Then, for training the in-context policy, we sample a short $h$-step trajectory randomly from dataset $\mathcal{D}^{i}$, and use the pretrained $E_{\phi}$ to transform raw transitions in this trajectory into compact task representations as
@@
Dynamic Programming. Classical RL algorithms learn a value function to backpropagate expected returns using dynamic programming updates, showcasing high stitching capacity that allows for not only imitating collected data but also achieving substantial improvement beyond it (Sutton et al., 1998). Inspired by this, we harness the stitching property to offer a promising avenue for unleashing ICRL’s potential toward explicit reward maximization. Following common practice in advanced offline RL algorithms (Kostrikov et al., 2022; Snell et al., 2023), we learn both a state value function $V$ and an action value function $Q$. As shown in Figure 1-(b), we design a prompt-based multi-head transformer architecture to simultaneously predict the optimal and value functions. Let $\theta$ denote parameters of the integrated transformer model. The model outputs the policy $\pi_{\theta}(a|s;\beta)$, the state value $V_{\theta}(s;\beta)$, and the action value $Q_{\theta}(s,a;\beta)$ using three separate heads, given the task prompt $\beta$ and corresponding query inputs ($s$ or $s,a$). Sec. 3.4 presents the detailed learning process.
@@ 3.3 World Model Pretraining for Prompt Construction
$\mathcal{L}(\phi,\varphi)=\mathbb{E}_{\eta_{t}^{i}\sim M^{i}}\!\left[\lVert[r_{t},s_{t+1}]-D_{\varphi}(s_{t},a_{t};z_{t}^{i})\rVert_{2}^{2}\mid z_{t}^{i}=E_{\phi}(\eta_{t}^{i})\right],~~~\forall~\text{task}~i.$ (3)
@@ 3.4 In-Context Q-Learning
where $L_{2}^{\omega}(u)=|\omega-\mathbbm{1}(u<0)|\cdot u^{2}$ is an asymmetric loss function with a expectile parameter $\omega\in(0.5,1)$. This form of expectile regression reduces the influence of Q<VQ\!<\!V predictions by a factor of $1\!-\!\omega$ while assigning more importance to Q>VQ\!>\!V predictions by a factor of $\omega$. In this way, we predict an upper-expectile of the temporal-difference target that approximates the maximum of $r(s_{t}^{i},a_{t}^{i})+\gamma Q_{\hat{\theta}}\left(s_{t+1}^{i},a_{t+1}^{i};\beta^{i}\right)$ over actions $a_{t+1}^{i}$ constrained to the dataset actions. More details can be found in Implicit Q-Learning (Kostrikov et al., 2022).
In-Context Policy Extraction. The value function learning procedure allows for stitching suboptimal trajectories to discover globally optimal behaviors. Then, we distill the in-context value functions into policy extraction with advantage-weighted regression (Peng et al., 2019), a supervised learning style that uses a simple and convergent maximum likelihood loss function as
@@ 3.5 Overall Optimization
$\mathcal{L}(\theta)=\mathsf{c}_{1}\,\mathcal{L}_{\pi}(\theta)+\mathsf{c}_{2}\,\mathcal{L}_{Q}(\theta)+\mathsf{c}_{3}\,\mathcal{L}_{V}(\theta),$ (7)
where we set coefficients $(\mathsf{c}_{1},\mathsf{c}_{2},\mathsf{c}_{3})$ to a balanced ratio of (1:1:1)(1\!:\!1\!:\!1) in all experiments. A detailed analysis of how different coefficient choices affect performance is provided in Appendix H. This design ensures that S-ICQL integrates lightweight prompt construction with a unified optimization pipeline, achieving efficient reward maximization and robust generalization across tasks.
@@ 4 Experiments
  • Can S-ICQL consistently outperform other strong baselines on unseen tasks? (Sec. 4.1)
@@
  • Does S-ICQL maintain strong performance on more complex and diverse environments? (Sec. 4.7)
@@
Baselines. We compare to six competitive ICRL approaches and one offline meta-RL method, including: 1) IC-IQL (Tarasov et al., 2025), 2) DIT (Dong et al., 2025), 3) DICP (Son et al., 2025), 4) IDT (Huang et al., 2024), 5) AD (Laskin et al., 2023), 6) DPT (Lee et al., 2023), and 7) UNICORN (Li et al., 2024b). Details about these baselines are provided in Appendix E.
@@ 4.1 Main Results
The baselines typically work under the few-shot setting, since they rely on task prompts or warm-start data to infer task representations. We evaluate S-ICQL against these baselines in an aligned few-shot setting, where all methods utilize the same number of interaction trajectories for task inference. Figure 2 illustrates the evaluation return curves across various environments using Mixed datasets, and Table 1 summarizes the numerical results of converged performance. Across environments with varying reward functions and transition dynamics, S-ICQL consistently demonstrates superior data efficiency and higher asymptotic performance. In more complex environments such as Ant-Dir and HalfCheetah-Vel, the advantage of S-ICQL is more pronounced, highlighting its promising in-context learning abilities on challenging tasks. Moreover, S-ICQL generally exhibits lower variance during the testing phase, indicating both improved data efficiency and better training stability. Notably, DPT yields suboptimal performance across most environments, underscoring the critical role of incorporating fundamental reward maximization into our framework. The significant improvement over DIT and IC-IQL also verifies that our method can provide a more efficient way for policy extraction with advantage-weighted regression or incorporating Q-learning objectives.
We also conduct an offline evaluation to assess the robustness of S-ICQL on trajectories collected by external behavior policies. Specifically, we use trajectories generated by non–S-ICQL policies and construct offline test sets by randomly sampling ten trajectories per task. We then evaluate DPT-style baselines on the same sampled trajectories to ensure a fair comparison. As shown in Table 2, S-ICQL maintains strong performance in the offline setting, indicating that its advantages persist even when interaction is restricted and evaluation relies solely on externally collected data.
Table 2: Offline evaluation returns of S-ICQL and baselines on Mixed datasets.
Methods Darkroom Push Reach Cheetah-Vel Walker-Param Ant-Dir
[…]
@@ 4.2 Ablation Study
Figure 3 shows the few-shot evaluation return curves of S-ICQL and its ablations on Mixed datasets across representative environments. Table 3 summarizes the numerical results of converged performance. First, ablating the world model causes a notable performance drop, especially in complex tasks like Ant-Dir, emphasizing its importance for precise in-context inference. Second, removing Q-learning reduces performance, underscoring its role in refining policies using suboptimal data. Integrating Q-learning enables S-ICQL to improve policies from noisy or inferior trajectories, enhancing its stitching ability to maximize rewards. Finally, removing both components results in the greatest decline, reducing the model to simple behavioral cloning that fails to generalize in complex tasks. Overall, the ablation results validate S-ICQL’s effectiveness in harnessing both world modeling and Q-learning for efficient reward maximization and generalization across tasks.
4.3 Generalization to Out-of-Distribution (OOD) Tasks
@@ 4.4 Stitching Capability
Figure 5: Comparison of best dataset returns with DPT, w/o_q, and S-ICQL on training tasks.
We conducted an in-depth analysis to examine whether S-ICQL truly exhibits stitching capability. Specifically, we pretrained the model on offline datasets containing only suboptimal trajectories and evaluated it on the corresponding training tasks. Figure 5 compares the best returns in the offline datasets (OPT) with the performance of DPT, w/o_q (S-ICQL without Q-learning), and S-ICQL, including both evaluation curves and final returns. As w/o_q and DPT follow the supervised pretraining paradigm, their performance never exceeds the dataset best, underscoring the limitations of purely supervised methods. In contrast, S-ICQL achieves performance exceeding the best in-dataset return on training tasks, providing strong evidence of genuine stitching by composing suboptimal trajectory segments into globally superior policies.